OpenStudy (anonymous):

What are the x– and y–intercepts of the graph of y = x^2 + x – 6?

OpenStudy (anonymous):

For finding x intercept, Put y = 0 And for finding y intercept, Put x = 0..

OpenStudy (anonymous):

So: for finding x intercept you will put y = 0 that is: \[x^2 + x - 6 = 0\] Try to factor this equation and you will get two values of x..

OpenStudy (anonymous):

Tell me what is your factors to this equation??

OpenStudy (anonymous):

*are

OpenStudy (anonymous):

to factor it i would fing two numbers that multiply to -6 and add to 0?

OpenStudy (anonymous):

Add to 1..

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Solving or not??

OpenStudy (anonymous):

-2 and 3 are my factors

OpenStudy (anonymous):

Then you are pretty correct in this..

OpenStudy (anonymous):

So is this right? \[(x+3)(x-2) = 0\]

OpenStudy (anonymous):

okay so those are my x and y factors?

OpenStudy (anonymous):

No these are only for x we will find y somewhat later..

OpenStudy (anonymous):

okay...can u show me the steps for finding y?

OpenStudy (anonymous):

For finding y, put x = 0 in your original equation..

OpenStudy (anonymous):

okay so y=0^2+0-6?

OpenStudy (anonymous):

Yes go ahead so what did you get for y ??

OpenStudy (anonymous):

-6?

OpenStudy (anonymous):

So true..

OpenStudy (anonymous):

thankyou so much!!! I was having alot of trouble with this!

OpenStudy (anonymous):

Welcome dear..